OpenStudy (anonymous):

does anyone know how to solve x+y=10

OpenStudy (anonymous):

u have only 1 equation and no other realtions??

OpenStudy (anonymous):

relations*

OpenStudy (anonymous):

no this is all i have

OpenStudy (anonymous):

oh yes...x+y=10 , y=x+8

OpenStudy (anonymous):

put y=x+8 in the first equation

OpenStudy (anonymous):

two variables,u must have two eqs to find the solution

OpenStudy (anonymous):

x+x+8=10 2x=2 x=1

OpenStudy (anonymous):

so from 2nd equation u get y=1+8=9

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

No

OpenStudy (anonymous):

where's the problem??

OpenStudy (anonymous):

your second equation says y=x+8. i put it in first equation so i get x+y=10 x+(x+8)=10 2x+8=10 2x=10-8 2x=2 x=2/2=1

OpenStudy (anonymous):

Got it...thank you

OpenStudy (anonymous):

you most welcome..........